State whether the following statements are true or false. Give reasons for your answers
(1) The function f:R^3®R, given by f( x, y, z) =|x|+|y|+|z| is differential at (2, 3,-1).
(2) The function f(x, y) =max{y/x, x} is a homogenous function on R^2.
(3) The domain of the f/g where f(x, y) =2xy and g(x, y) =x^2+y^2 is R^2.
1
Expert's answer
2020-03-16T12:08:42-0400

1)It is true. In "B_1(2,3,-1)=\{(x,y,z)\mid |(2,3,-1)-(x,y,z)|<1\}" we have "f(x,y,z)=x+y-z". It is a differentiable in "B_1(2,3,-1)" function, so it is differentiable in "(2,3,-1)".

2)It is not true. For "x_0=1,y_0=1,\lambda_0=0" we have "\lambda_0^kf(x_0,y_0)=0^k\max\{\frac{1}{1},1\}=1" is defined, but "f(\lambda_0x_0,\lambda_0y_0)=\max\{\frac{\lambda_0y_0}{\lambda_0x_0},\lambda_0x_0\}=\max\{\frac{0}{0},0\}" is not defined, because "\frac{0}{0}" is not defined, so "\lambda_0^kf(x_0,y_0)\neq f(\lambda_0x_0,\lambda_0y_0)".

3)It is false, beacuse "\frac{f}{g}" is not continuous at "(0,0)", so we cannot define "\frac{f}{g}" at this point.

"\frac{f}{g}" is not continuous at "(0,0)", because there is not a "\lim\limits_{(x,y)\to (0,0)} \frac{f(x,y)}{g(x,y)}". Indeed, for "\varepsilon=1" we have that for every "\delta>0" we can take "\left(0,\frac{\delta}{2}\right),\left(\frac{\delta}{2},\frac{\delta}{2}\right)\in B_{\delta}(0)=\{(x,y)\mid |(x,y)|<\delta\}" and "\left|\frac{f\left(0,\frac{\delta}{2}\right)}{g\left(0,\frac{\delta}{2}\right)}-\frac{f\left(\frac{\delta}{2},\frac{\delta}{2}\right)}{g\left(\frac{\delta}{2},\frac{\delta}{2}\right)}\right|=|0-1|\ge\varepsilon", so "\frac{f}{g}" is not satisfy the Cauchy test of limit existence at "(0,0)".